OpenStudy (anonymous):

I have to solve in Algebra form: Sam can row 10 miles downstream in one third the time he can row 6 miles upstream. If Sam can row 5 miles per hour in still water, how fast is the current?

OpenStudy (anonymous):

Let T be the time Sam takes to row 6 miles upstream and C the current speed. t = d / r \[t=\frac{6}{5-c}\] is the expression for the 6 mile upstream time and\[\frac{t}{3}= \frac{10}{5+c}\] is the equation for the 10 mile downstream time. Solve both for c and t.\[c= 3\frac{1}{3} \text{mph} \text{ and } t=\frac{18}{5}\text{hours}\text{ or 3}\text{ hrs } 36 \text{ minutes} \]